What happens to the centripetal acceleration of a revolving body if you double the orbital speed v and half the angular velocity ? Explain?

Grade:11

2 Answers

prateek
7 years ago
centripetal accleration = linear velocity^2 \radius
radius x angular velocity = linear velocity
then
V=2v
R=r\4
putting in formula we get
acceleration increases by 8 times
Pratyush Ranjan Roul
7 years ago
v2=2v  w2=w/2
but r = v/w,
r2=v/w2 =2v/(w/2)= 4v/w = 4r
a2=(v2)2/r2
a2=4v2/ 4r
So, acceleration is unchanged